# A Survey for Deep Reinforcement Learning Based Network Intrusion Detection

### v1

@@ Abstract
Cyber-attacks are gradually becoming more sophisticated and highly frequent nowadays, and the significance of network intrusion detection systems have become more pronounced. This paper investigates the prospects and challenges of employing deep reinforcement learning technologies in network intrusion detection. It begins with an introduction to the fundamental theories and technological frameworks of deep reinforcement learning including classic deep Q-network and actor-critic algorithms, followed by a review of essential research that has leveraged deep reinforcement learning for network intrusion detection in recent years. This research assesses these challenges and efforts in terms of model training efficiency, the detection capabilities for minority and unknown class attacks, improved network feature selection and unbalanced dataset issues. Performances of deep reinforcement learning models are comprehensively investigated. The findings reveal that although deep reinforcement learning shows promise in network intrusion detection, many of the latest deep reinforcement learning technologies are yet to be fully explored. Some deep reinforcement learning based models can achieve state-of-the-art results in some public datasets, in some cases, even better than traditional deep learning methods. The paper concludes with recommendations for the enhanced deployment and testing of deep reinforcement learning technologies in real-world network scenarios to further improve their application. Special emphasis is placed on the Internet of Things intrusion detection. We offer discussions on recently proposed deep architectures, revealing possible future policy functions used for deep reinforcement learning based network intrusion detection. In the end, we propose integrating deep reinforcement learning and broader generative methods and models to assist and further improve their performance. These advancements aim to address the current gaps and facilitate more robust and adaptive network intrusion detection systems.
Index Terms:
Intrusion Detection, Deep Reinforcement Learning, Cyber-Security, Cyber-Physical Systems
@@ I Introduction
Cybersecurity is a significant challenge in the age of global informatization [1]. As digital lifestyles become increasingly prevalent, our reliance on cybersecurity and the need for itare growing rapidly as well [2]. Nowadays, cybercriminal activities are inflicting substantial economic losses on various industrial and government sectors. According to live data from the AV-TEST Institute, more than 450,000 new malicious programs are detected every day, with over 1.2 billion instances of malware emerging within 2023 alone, and these numbers are still on the rise [3]. These malicious programs infiltrate personal and corporation digital systems through a wide range of vulnerabilities, posing serious threats to personal privacy, commercial secrets, and other sensitive records. It is estimated that cybersecurity-related issues have caused a total loss of up to 400 billion US dollars to the global economy [4]. From the introduction of cybersecurity ventures, cybercrime has caused \$1.5 trillion in losses [5] in 2019, a figure that is expected to climb to \$9.5 trillion US dollars by 2024 [6]. Moreover, critical national infrastructures from energy, healthcare sectors, and port automation systems, are becoming primary targets for cyber-attackers [7]. According to 2024 World Economic Forum, 94% of government leaders and business executives believe that their organizations are still at a lower defence level when facing cyber-attacks [8]. In short, the importance of cybersecurity is set to grow significantly worldwide.
@@
Network Intrusion Detection (NID) is a crucial defence mechanism in the field of cybersecurity. It effectively protects computers and other digital devices from external attacks [9]. It was first proposed in 1994 [10] and later described as integrating information extracted from computers to identify resource abuse within the network and attacks originating from outside entities [11]. Basically, intrusion detection systems (IDS) can be categorized into Network Intrusion Detection Systems (NIDS), which is based on the observation of network traffic between different nodes [12] and Host-based Intrusion Detection Systems (HIDS), which means monitoring activities on a specific host, including applications being used and file systems being accessed [13]. The primary purpose of NID is to prevent network attacks by identifying abnormal traffic or access operations [9]. Objectives of network attacks are becoming increasingly complex, traditional signature-based methods that identifying known attacks based on pattern matching of known signatures have fallen behind anomaly-based detection approaches [14]. Since the anomaly-based detection has a higher efficiency and dynamic adaptability, it is now widely accepted by the NID community [15].
@@
Figure 1: An overview of network intrusion detection and specific working flow of anomaly-based netwotk intrusion detection.
@@
Artificial intelligence (AI) plays an essential role in NID [7, 16]. Traditional machine learning (ML) algorithms, including supervised and unsupervised learning e.g., Support Vector Machines (SVM), K-Nearest Neighbors (KNN), Random Forests (RF), and Multilayer Perceptron (MLP), have made improvements for NIDS in multiple ways [17]. Further, deep learning (DL), by constructing deep neural networks, are capable of learning and fitting highly complex patterns and features from large amounts of given training data [18]. It can effectively learn and simulate patterns of normal behaviour in network traffic, thereby identifying abnormal activities or potential intrusions [16, 17, 18]. Compared with traditional ML methods, critical features in network intrusion could be extracted automatically by deep learning without much laborious feature selection.
@@
Some surveys focused on the general AI in intrusion detection [33] or DRL in general cybersecurity [21]. However, there has been limited research on applying deep reinforcement learning to network intrusion detection. Thus, a need for a more detailed and comprehensive survey specifically focused on the DRL and network intrusion detection. This paper primarily concentrates on the exploration of DRL applications within the domain of NID over the past five years. It aims to provide a systematic review of the most current advancements in RL applications for NID, endeavoring to elucidate how RL is revolutionizing and enhancing NID systems. To achieve this, firstly, preliminary knowledge for RL is presented including the Markov decision process, Q-learning, Deep Reinforcement Learning, Inverse Reinforcement Learning and how we evaluate the performance of RL-based NID model. Then, we retrospectively investigate representative works of DRL-based NID in the past few years focused on the overview of the most often used datasets, efforts on network feature engineering, handling unbalanced datasets, improving training efficiency and identifying minority intrusions. Finally, we present our general discussion based on these representative works hoping to provide some ideas and future research directions.
Figure 2: Number of related publications start from 2017 (Taking reinforcement learning and network intrusion as searching key words, data is collected from Web of Science).
@@ II-A Overview
Classic RL system is, in general, comprised of an environment, agent, policy, reward, and value function. [34]. Agent means a decision-making, goal-seeking and highly interactive virtual entity. Environment refers to everything that the agent interact with: applying action to it and receiving feedback from it. Policy $\pi$ determines which action the agent takes when in a given state of the environment. It could be a function or a simple lookup matrix. Reward is quantitative feedback from the environment after the agent takes one specific action followed by a state. The positive and negative reward describes how “good” or “bad” the action is regarding to the final goal of the agent. Value function is used to evaluate the quality level of a state or action. Thus, it is divided into state-value function and action-value function.
@@ II-B Markov decision process
Markov decision process (MDP) is the principal framework for decision in stochastic and uncertain environment [35]. It assumes that an agent can observe the current state $s_{t}$, and choose to take an action $a_{t}$. Then, the agent moves to the next state $s_{t+1}$. Normally, it is described as a tuple $\left(S,A,P,R,\gamma\right)$ with the following five essential elements. $S$ is a state space, including all states that can be observed by the agent. $A$ is an action space, i.e., the set of all actions that can be taken by the agent. $p\left(s^{{}^{\prime}}|s,a\right)$ describes the probability of transferring to a specific state $s^{{}^{\prime}}$ by taking action $a$ in the given state $s$. Reward function $r\left(\right.s,a,s^{{}^{\prime}}\left.\right)$ gives the reward (positive or negative) that the agent get when making a transition from state $s$ to $s^{{}^{\prime}}$ by taking action $a$. $\gamma$ is a discount factor that can be used to indicate short-term or long-term importance of rewards. The fundamental property of an MDP is that the next state $s_{t+1}$ depends only on the current state $a_{t}$ and the action $a_{t}$ that the agent has taken. The interaction process between the agent and the environment in a Markov decision process is well illustrated in Figure 3.
Figure 3: Agent interaction with environment based MDP in RL.
The final goal of the agent is to maximize long-term rewards. State-value function $V\left(\right.s\left.\right)$ can estimate the value of a state, which starts from a specific state $s$ following the action chosen by current policy $\pi$ of the agent. It can be utilized to approximate how much long-term reward the agent can gain in all next states based on the current $\pi$ shown in Equation 1.
@@
Firstly, the agent follows policy $\pi$ to take an action $a$ in state $s$ to gain an immediate reward $R(s,a)$. And over all possible next states $s^{\prime}$, we have discounted sum of possible $V(s^{\prime})$ by adding discounted factor $\gamma$. Furthermore, Bellman optimality equations define optimal state-value function $V^{*}(s)$ to get the maximum return for each state and optimal action-value function $Q^{*}(s,a)$ to get the maximum return for each pair of state-action, shown in Equation 2 and Equation 3.
@@
$Q^{*}(s,a)=R(s,a)+\gamma\sum_{s^{\prime}\in S}p(s^{\prime}|s,a)V^{*}(s^{\prime})$ (3)
@@ II-C Q-learning
Q-learning is a classic value-based algorithm in RL [36]. It allows the agent to learn how to select actions in a given state to maximize the expected total reward. [37]. The agent updates the value of action-state pair (Q value) by exploring possible actions of the state ($\epsilon$-greedy strategy [38]). The upgrade of Q value is based on the current Q value, immediate reward, and maximum Q value of next state. Equation 4 is normally used to update the Q value. In the equation, the Q value $Q(s_{t},a_{t})$ represents the expected utility of taking action $a_{t}$ in the current state $s_{t}$, considering both the immediate reward $r_{t}$ and the expected rewards of future states. The immediate reward $r_{t}$ is the gain obtained from taking action $a_{t}$.
@@
In the current state $s_{t}$, the agent has a high sampling probability to take action $a_{t}$ as $Q(s_{t},a_{t})$ holds the highest value among all actions that could be taken in $s_{t}$. [^1] The agent still has small chances to choose other actions although they have a relatively lower Q value $r_{t}$ is the immediate reward that the agent obtained by taking $a_{t}$. $Q(s_{t+1},a)$ means all possible Q value corresponding to all possible action $a$ in $s_{t+1}$. Discounted factor $\gamma$ is used to determine if the agent should focus on long-term reward, the smaller $\gamma$ is, the shorter-sighted it is. In the end, learning rate $\alpha$ determines how much fresh learning experience should be added to past experience.
@@ II-D Deep Reinforcement Learning
Figure 4: Classic deep Q network learning process
@@ II-D1 Deep Q Network, DQN
Deep reinforcement learning (DRL) is a major progress in RL [39]. It makes agent could input high-dimensional data e.g. images and audio, which are difficult for typical Q-learning since they may need a massive tabular to store the value of state-action pair. It is normally unacceptable for memory. DQN [40] is a classic DRL algorithm that leverages the dynamic transition of experience replay and utilization of a target network.
@@
Specifically, the agent interacts with the environment using random strategy and initial hyper-parameters $\theta_{0}$ to get some initial experience $(s_{t},a_{t},r_{t},s_{t+1})$ to fill the buffer of experience replay. Based on this initial knowledge, the target network (also initialized by $\theta_{0}$) could calculates the target Q value by using the Bellman equation for data in replay buffer. Then, in the training process, a minibatch of experience is randomly chosen from experience replay buffer to update the hyper-parameters of the policy network $\pi_{\theta_{0}}$ (Q-Network), by minimum the loss of predicted Q value $Q(s_{t},a;\theta)$ and target Q value, shown in the Equation 5
$L(\theta)=E[(r_{t}+\gamma\max_{a}Q(s_{t+1},a;\theta^{\prime})-Q(s_{t},a;\theta))^{2}]$ (5)
After the policy network is updated into $\theta^{\prime}$, it starts to interact with the environment again to accumulate new experiences, which will be used to partially update experiences replay buffer. And the hyper-parameter of the target network will be frozen as $\theta_{0}$ and subsequently updated into $\theta^{\prime}$ after a fixed time steps. Parameter upgrades of target network always lag behind the upgrades of the policy network. The training will stop when the maximum accumulated reward has been achieved or the loss has converged to a stable range.
@@ II-D2 Policy Gradient
The policy gradient method [41] is employed to optimize decision-making policies in reinforcement learning directly, without relying on a value function. The principle, different from value function-based methods, is to optimize policy parameters directly and then to achieve the maximum total reward. Policy Gradient Theorem is the very foundation of policy gradient. Let a parametric policy $\pi(a|s,\theta)$, choosing action $a$ given a state $s$ by a probability $p=\pi_{\theta}$, then, the performance of the policy can be quantified as follows. $R(\tau)$ means return from a single trace. And policy gradient can be described in Equation 6, Equation 7 and Equation 8:
@@
$\nabla_{\theta}J(\theta)=E_{s,a\sim\pi_{\theta}}\left[\sum_{t=0}^{T}\nabla_{\theta}\log\pi_{\theta}(a_{t}|s_{t})G_{t}\right]$ (7)
@@
where $k$ represents a future time point starting from a specific initial time denoted by $t$. As time passes by, $k$ gradually increases. It is desired that the contribution of rewards received at farther future time points to the current state decreases gradually, hence $\gamma$, the discount factor, is reduced over time. The initial trajectories are gathered from random interactions between the agent and the environment over a fixed duration. These trajectories are used only once. After all trajectories in a set have been utilized, the agent restarts interactions with the environment to collect a new set of trajectories for the next round of training. However, when having large gradient updates in training, the process becomes unstable. To address this, researchers proposed Trust Region Policy Optimization (TRPO) [42] and Proximal Policy Optimization (PPO) [43] to further improve the stability of training.

### v2

@@ Abstract
Cyber-attacks are gradually becoming more sophisticated and highly frequent nowadays, and the significance of network intrusion detection systems has become more pronounced. This paper investigates the prospects and challenges of employing deep reinforcement learning technologies in network intrusion detection. It begins with an introduction to the fundamental theories and technological frameworks of deep reinforcement learning including classic deep Q-network and actor-critic algorithms, followed by a review of essential research that has leveraged deep reinforcement learning for network intrusion detection in recent years. This research assesses these challenges and efforts in terms of model training efficiency, the detection capabilities for minority and unknown class attacks, improved network feature selection and unbalanced dataset issues. Performances of deep reinforcement learning models are comprehensively investigated. The findings reveal that although deep reinforcement learning shows promise in network intrusion detection, many of the latest deep reinforcement learning technologies are yet to be fully explored. Some deep reinforcement learning based models can achieve state-of-the-art results in some public datasets, in some cases, even better than traditional deep learning methods. The paper concludes with recommendations for the enhanced deployment and testing of deep reinforcement learning technologies in real-world network scenarios to further improve their application. Special emphasis is placed on the Internet of Things intrusion detection. We offer discussions on recently proposed deep architectures, revealing possible future policy functions used for deep reinforcement learning based network intrusion detection. In the end, we propose integrating deep reinforcement learning and broader generative methods and models to assist and further improve their performance. These advancements aim to address the current gaps and facilitate more robust and adaptive network intrusion detection systems.
@@ I Introduction
Cybersecurity is a significant challenge in the age of global informatization [4]. As digital lifestyles become increasingly prevalent, our reliance on cybersecurity and the need for it is growing rapidly as well [106]. Nowadays, cybercriminal activities are inflicting substantial economic losses on various industrial and government sectors. According to live data from the AV-TEST Institute, more than 450,000 new malicious programs are detected every day, with over 1.2 billion instances of malware emerging within 2023 alone, and these numbers are still on the rise [30]. These malicious programs infiltrate personal and corporation digital systems through a wide range of vulnerabilities, posing serious threats to personal privacy, commercial secrets, and other sensitive records. It is estimated that cybersecurity-related issues have caused a total loss of up to 400 billion US dollars to the global economy [28]. According to Cybersecurity Ventures, cybercrime has caused \$1.5 trillion in losses [136] in 2019, a figure that is expected to climb to \$9.5 trillion US dollars by 2024 [137]. Moreover, critical national infrastructures from energy, healthcare sectors, and port automation systems, are becoming primary targets for cyber-attackers [27]. According to 2024 World Economic Forum, 94% of government leaders and business executives believe that their organizations are still at a lower defence level when facing cyber-attacks [18]. In short, the importance of cybersecurity is set to grow significantly worldwide.
@@
Network Intrusion Detection (NID) is a crucial defence mechanism in the field of cybersecurity. It effectively protects computers and other digital devices from external attacks [131]. It was first proposed in 1994 [81] and later described as integrating information extracted from computers to identify resource abuse within the network and attacks originating from outside entities [12]. Basically, intrusion detection systems (IDS) can be categorized into Network Intrusion Detection Systems (NIDS), which is based on the observation of network traffic between different nodes [129] and Host-based Intrusion Detection Systems (HIDS), which means monitoring activities on a specific host, including applications being used and file systems being accessed [88]. The primary purpose of NID is to prevent network attacks by identifying abnormal traffic or access operations [131]. Objectives of network attacks are becoming increasingly complex, traditional signature-based methods that identify known attacks based on pattern matching of known signatures have fallen behind anomaly-based detection approaches [17]. Since the anomaly-based detection has a higher efficiency and dynamic adaptability, it is now widely accepted by the NID community [10].
@@
Figure 1: An overview of network intrusion detection and specific working flow of anomaly-based network intrusion detection.
@@
Artificial intelligence (AI) plays an essential role in NID [27, 23]. Figure 1 shows an overview of network intrusion detection and specific working flow of anomaly-based network intrusion detection. The system is positioned between the Internet and internal users, with traffic flowing through a router and firewall before reaching the Network Intrusion Detection System, which employs both signature-based detection for known attack patterns and anomaly-based detection for novel threats. The anomaly-based detection pipeline, detailed in the lower section, consists of four sequential stages: network traffic collection from network devices, data processing for feature extraction and normalization, automated model prediction using neural network algorithms, and alert raising when potential intrusions are detected. This dual-detection approach combines traditional pattern-matching techniques with advanced machine learning methods to provide robust, real-time protection against both known and emerging network security threats. Traditional machine learning (ML) algorithms, including supervised and unsupervised learning e.g., Support Vector Machines (SVM), K-Nearest Neighbors (KNN), Random Forests (RF), and Multilayer Perceptron (MLP), have made improvements for NIDS in multiple ways [3]. Further, deep learning (DL), by constructing deep neural networks, is capable of learning and fitting highly complex patterns and features from large amounts of given training data [65]. It can effectively learn and simulate patterns of normal behaviour in network traffic, thereby identifying abnormal activities or potential intrusions [23, 3, 65]. Compared with traditional ML methods, critical features in network intrusion could be extracted automatically by deep learning without as much laborious feature selection.
@@
Some surveys focused on the general AI in intrusion detection [122] or DRL in general cybersecurity [92]. However, there has been limited research on applying deep reinforcement learning to network intrusion detection. Thus, there is a need for a more detailed and comprehensive survey specifically focused on DRL and network intrusion detection. This paper primarily concentrates on the exploration of DRL applications within the domain of NID over the past five years. It aims to provide a systematic review of the most current advancements in RL applications for NID, endeavoring to elucidate how RL is revolutionizing and enhancing NID systems. To achieve this, firstly, preliminary knowledge of RL is presented including the Markov decision process, Q-learning, Deep Reinforcement Learning, Inverse Reinforcement Learning and how we evaluate the performance of RL-based NID model. Then, we retrospectively investigate representative works of DRL-based NID in the past few years focused on the overview of the most often used datasets, efforts on network feature engineering, handling unbalanced datasets, improving training efficiency and identifying minority intrusions. Finally, we present our general discussion based on these representative works hoping to provide some ideas and future research directions.
Figure 2: Number of related publications since 2017, based on a Web of Science search using the keywords “reinforcement learning” and “network intrusion.”
@@ II-A Overview
A classic RL system is, in general, comprised of an environment, agent, policy, reward, and value function. [54]. Agent means a decision-making, goal-seeking and highly interactive virtual entity. Environment refers to everything that the agent interacts with: applying action to it and receiving feedback from it. Policy $\pi$ determines which action the agent takes when in a given state of the environment. It could be a function or a simple lookup matrix. Reward is quantitative feedback from the environment after the agent takes one specific action followed by a state. The positive and negative reward describes how “good” or “bad” the action with regarding to the final goal of the agent. Value function is used to evaluate the quality level of a state or action. Thus, it is divided into state-value function and action-value function.
@@ II-B Markov decision process
Markov decision process (MDP) is the principal framework for decision in stochastic and uncertain environment [98]. It assumes that an agent can observe the current state $s_{t}$, and choose to take an action $a_{t}$. Then, the agent moves to the next state $s_{t+1}$. Normally, it is described as a tuple $\left(S,A,P,R,\gamma\right)$ with the following five essential elements. $S$ is a state space, including all states that can be observed by the agent. $A$ is an action space, i.e., the set of all actions that can be taken by the agent. $p\left(s^{{}^{\prime}}|s,a\right)\in P$ describes the probability of transferring to a specific state $s^{{}^{\prime}}$ by taking action $a$ in the given state $s$. Reward function $R:S\times A\rightarrow\mathbb{R}$ gives the expected immediate reward $R(s,a)=\mathbb{E}[r_{t}|s_{t}=s,a_{t}=a]$ that the agent receives after taking action $a$ in state $s$. Alternatively, the reward can be defined as $r(s,a,s^{\prime})\in\mathbb{R}$, representing the reward received when transitioning from state $s$ to $s^{\prime}$ via action $a$, where $R(s,a)=\sum_{s^{\prime}\in S}p(s^{\prime}|s,a)r(s,a,s^{\prime})$. $\gamma$ is a discount factor that can be used to indicate short-term or long-term importance of rewards. The fundamental property of an MDP is that the next state $s_{t+1}$ depends only on the state $s_{t}$ and the action $a_{t}$ that the agent has taken. The interaction process between the agent and the environment in a Markov decision process is well illustrated in Figure 3.
Figure 3: Agent interaction with environment based on MDP in RL.
[…]
@@
Firstly, the agent follows policy $\pi$ (decision-making policy) to take an action $a$ in state $s$ to gain an immediate reward $R(s,a)$. And over all possible next states $s^{\prime}$, we have discounted sum of possible $V(s^{\prime})$ (expected future returns) by adding discount factor $\gamma$ (future reward weight). Furthermore, Bellman optimality [103] equations define optimal state-value function $V^{*}(s)$ to get the maximum return for each state and optimal action-value function $Q^{*}(s,a)$ to get the maximum return for each state-action pair, shown in Equation 2 and Equation 3.
@@
$Q^{*}(s,a)=R(s,a)+\gamma\sum_{s^{\prime}\in S}p(s^{\prime}|s,a)\max_{a^{\prime}}Q^{*}(s^{\prime},a^{\prime})$ (3)
@@ II-C Q-learning
Q-learning is a classic value-based algorithm in RL [141]. It allows the agent to learn how to select actions in a given state to maximize the expected total reward. [19]. The agent updates the value of action-state pair (Q value) by exploring possible actions of the state ($\epsilon$-greedy strategy [118]). The update of Q value is based on the current Q value, immediate reward, and maximum Q value of next state. Equation 4 is normally used to update the Q value. In the equation, the Q value $Q(s_{t},a_{t})$ represents the expected utility of taking action $a_{t}$ in the current state $s_{t}$, considering both the immediate reward $r_{t}$ and the expected rewards of future states.
@@
Q-learning typically employs an $\epsilon$-greedy exploration strategy: with probability $1-\epsilon$, the agent selects the action $a_{t}=\arg\max_{a}Q(s_{t},a)$ that maximizes the Q-value (exploitation); with probability $\epsilon$, it randomly selects an action from the action space (exploration). This balances exploration of the environment with exploitation of learned knowledge. $r_{t}$ is the immediate reward that the agent obtained by taking $a_{t}$. $Q(s_{t+1},a)$ means all possible Q value corresponding to all possible action $a$ in $s_{t+1}$. Discounted factor $\gamma$ is used to determine if the agent should focus on long-term reward, the smaller $\gamma$ is, the more short-sighted the agent becomes. Finally, the learning rate $\alpha\in(0,1]$ controls the magnitude of the Q-value update, determining the extent to which newly observed information overrides the existing estimate. A higher $\alpha$ places more weight on recent experience, while a lower $\alpha$ preserves more historical knowledge.
@@ II-D Deep Reinforcement Learning
Figure 4: Classic deep Q network process
@@ II-D1 Deep Q Network, DQN
Deep reinforcement learning (DRL), shown in Figure 4, is a major progress in RL [58]. It allows agents to handle high-dimensional inputs (e.g., images and audio), which pose challenges for typical Q-learning as they would require a massive table to store state-action pair values—typically unacceptable given memory constraints. DQN [41] is a classic DRL algorithm that leverages the dynamic transition of experience replay and utilization of a target network.
@@
Specifically, the agent interacts with the environment using an $\epsilon$-greedy strategy (typically with high initial $\epsilon$ for exploration) and network parameters initialized as $\theta_{0}$ to collect initial experiences $(s_{t},a_{t},r_{t},s_{t+1})$ that populate the experience replay buffer. Based on this initial knowledge, the target network (also initialized by $\theta_{0}$) could calculates the target Q value by using the Bellman equation for data in replay buffer. Then, in the training process, a minibatch of transitions is randomly sampled from the experience replay buffer to update the parameters $\theta$ of the Q-network by minimizing the loss function shown in Equation 5. The loss measures the mean squared error between the predicted Q-value $Q(s_{t},a_{t};\theta)$ and the target Q-value $r_{t}+\gamma\max_{a^{\prime}}Q(s_{t+1},a^{\prime};\theta^{-})$, where $\theta^{-}$ denotes the parameters of the target network.
$L(\theta)=\mathbb{E}[(r_{t}+\gamma\max_{a^{\prime}}Q(s_{t+1},a^{\prime};\theta^{-})-Q(s_{t},a_{t};\theta))^{2}]$ (5)
After the Q-network parameters are updated to $\theta$, the agent continues to interact with the environment to collect new experiences, which are added to the experience replay buffer (oldest experiences may be discarded if the buffer is full). The target network parameters $\theta^{-}$ remain fixed for a predetermined number of training steps (e.g., every $C$ steps) and are then periodically synchronized with the current Q-network parameters: $\theta^{-}\leftarrow\theta$. This periodic update mechanism ensures that the target values remain stable during training, which is crucial for convergence. Alternatively, soft updates can be used: $\theta^{-}\leftarrow\tau\theta+(1-\tau)\theta^{-}$ where $\tau\ll 1$. The training will stop when the maximum accumulated reward has been achieved or the loss has converged to a stable range.
@@ II-D2 Policy Gradient
The policy gradient method [115] is employed to optimize decision-making policies in reinforcement learning directly, without relying on a value function. The principle, different from value function-based methods, is to optimize policy parameters directly and then to achieve the maximum total reward. Policy Gradient Theorem [126] is the very foundation of policy gradient. Let a parametric policy $\pi(a|s,\theta)$, choosing action $a$ given a state $s$ by a probability $p=\pi_{\theta}$, then, the performance of the policy can be quantified as follows. $R(\tau)$ means return from a single trace. It refers to one complete trajectory or episode of sequential interactions between an agent and environment, consisting of a sequence of states, actions, and rewards from start to finish. A policy gradient can be described in Equation 6, Equation 7 and Equation 8:
@@
$\nabla_{\theta}J(\theta)=\mathbb{E}_{\tau\sim\pi_{\theta}}\left[\sum_{t=0}^{T-1}\nabla_{\theta}\log\pi_{\theta}(a_{t}|s_{t})G_{t}\right]$ (7)
@@
where $T$ denotes the time horizon of an episode (which may be finite or infinite depending on the problem setting), and $\tau=(s_{0},a_{0},r_{0},s_{1},a_{1},r_{1},...,s_{T})$ represents a complete trajectory, and $k$ represents a future time point starting from a specific initial time denoted by $t$. As time passes by, $k$ gradually increases. It is desired that the contribution of rewards received at farther future time points to the current state decreases gradually. This is achieved through the exponential term $\gamma^{k-t}$, where $\gamma\in[0,1]$ is a fixed discount factor that remains constant throughout training. The geometric decay $\gamma^{k-t}$ ensures that rewards further in the future have progressively smaller influence on the current state value. The initial trajectories are gathered from random interactions between the agent and the environment over a fixed duration. These trajectories are used only once. After all trajectories in a set have been utilized, the agent restarts interactions with the environment to collect a new set of trajectories for the next round of training. However, when having large gradient updates in training, the process becomes unstable. To address this, researchers proposed Trust Region Policy Optimization (TRPO) [107] and Proximal Policy Optimization (PPO) [108] to further improve the stability of training.